OpenStudy (anonymous):

the arch of an underpass is a semi ellipse 60 feet wide and 20 feet high, find the clearance at the edge of a lane if the edge is 20 ft at the middle? my answer was 25ft. am i right??

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (anonymous):

getting the c..

OpenStudy (anonymous):

i don't know.. i dont understand the clearance at the edge of a lane..

OpenStudy (anonymous):

First, we need to get the axis.

OpenStudy (anonymous):

The minor axis is \(20\) and the semi-major axis is \(60/2=30\)

OpenStudy (anonymous):

Our equation is: \[ \frac{x^2}{30^2}+\frac{y^2}{20^2}=1 \]

OpenStudy (anonymous):

"the edge is 20 ft at the middle" This means \(x=20\)

OpenStudy (anonymous):

So solve for \(y\): \[ \frac{20^2}{30^2}+\frac{y^2}{20^2}=1 \]

OpenStudy (anonymous):

\[ \frac{y^2}{400}=1-\frac 49 \]

OpenStudy (anonymous):

\[ y^2=400\cdot \frac 59 \]

OpenStudy (anonymous):

15ft

OpenStudy (anonymous):

tsk

OpenStudy (anonymous):

\[ y=\pm \frac{20}{3}\sqrt{5} \]

OpenStudy (anonymous):

aaaaaaaah so heartbreaking..

OpenStudy (anonymous):

We can disregard the \(-\) since we know it is positive.

OpenStudy (anonymous):

thank you thank you

OpenStudy (anonymous):

The answer is \(\approx 15\) if you don't mind rounding.

OpenStudy (anonymous):

The exact answer is: \[ \frac{20\sqrt 5}{3} \]

OpenStudy (anonymous):

so? does that makes sense?

OpenStudy (anonymous):

Yeah.